OpenStudy (anonymous):

Use the “difference of squares” rule to factor the following expression: 9x^4-25y^2

OpenStudy (anonymous):

Oaky difference of square: a^2-b^2=(a+b)(a-b) So all we have to do is rewrite this in such a way that we have it in the form above^ (3x^2)^2-(5y)^2=(3x^2+5y)(3x^2-5y)

OpenStudy (anonymous):

okay so that formula is used for this kind of problem okay